OpenStudy (aryana_maria2323):

The particular solution of the differential equation dy dt equals y over 4 for which y(0) = 20 is a. y = 20e−0.25t b. y = 19 + e0.25t c. y = 20 e0.25t d. y = 20e4t

OpenStudy (aryana_maria2323):

@across

OpenStudy (aryana_maria2323):

Hello @across

OpenStudy (aryana_maria2323):

Can you help @ganeshie8

ganeshie8 (ganeshie8):

familiar with separation of variables ?

OpenStudy (aryana_maria2323):

No I just learned this subject today and I am not quite understanding it.

OpenStudy (aryana_maria2323):

Okay.

OpenStudy (aryana_maria2323):

Okay so that kindof cleared it up for me but I am still a little bit fuzzy in some places.

ganeshie8 (ganeshie8):

They key thing here is to get get x terms and dx on one side y terms and dy on the other side

ganeshie8 (ganeshie8):

x is t in our case

OpenStudy (aryana_maria2323):

Okay I got that part. What would be my first step in solving the problem?

ganeshie8 (ganeshie8):

Could you show me what you get after separating ?

OpenStudy (aryana_maria2323):

\[\frac{ dy }{ dt }= \frac{ y }{ 4 } \] after separated \[ydy=4dt\] Is that correct?

ganeshie8 (ganeshie8):

Nope, but its a good try. Try again...

OpenStudy (aryana_maria2323):

What am I doing wrong? because in the video he multiplied both sides with y and dy or in this case dt

ganeshie8 (ganeshie8):

The goal is to get dy and y terms on one side

OpenStudy (aryana_maria2323):

Ohh not on opposite sides like in the video?

ganeshie8 (ganeshie8):

give it a try

OpenStudy (aryana_maria2323):

ydydt=4

OpenStudy (aryana_maria2323):

I mean ydy-dt=4

ganeshie8 (ganeshie8):

Heyy its just algebra : \[\frac{ dy }{ dt }= \frac{ y }{ 4 }\] divide \(y\) both sides and get \[\dfrac{1}{y}\dfrac{dy}{dt} = \dfrac{1}{4}\] multiply \(dt\) both sides and get \[\dfrac{1}{y}dy = \dfrac{1}{4}dt\]

OpenStudy (aryana_maria2323):

Okay why do you have the 1 over the y and the 4?

ganeshie8 (ganeshie8):

that is a good question

OpenStudy (aryana_maria2323):

I thought you said that we should have the dy and y terms on one side?

OpenStudy (aryana_maria2323):

Can you please show me the steps and walk me through them so I can understand how to do the problem?

ganeshie8 (ganeshie8):

Consider below equation : \[x=2y\]

OpenStudy (aryana_maria2323):

Okay

ganeshie8 (ganeshie8):

divide \(2\) both sides, what do you get ?

OpenStudy (aryana_maria2323):

\[\frac{ dy }{ dx }=\frac{ 1 }{ 2 }\] \[\frac{ dx }{ dy }=2\]

OpenStudy (aryana_maria2323):

@mathmale @across @jim_thompson5910 @retirEEd

OpenStudy (across):

\[\ln y=\frac14x+c\\y=ce^{\frac x4}\\20=c\\y=20e^{\frac x4}\]

OpenStudy (aryana_maria2323):

so is that C or D? @across

OpenStudy (mathmale):

dy dt equals y over 4:\[\frac{ dy }{ dt}=\frac{ y }{ 4 }\]...after cross-mult., becomes (dy/y)=(dt/4), or (dy/y)=(1/4)dt.

OpenStudy (aryana_maria2323):

Thank you I got it correct.

OpenStudy (mathmale):

Glad you got it right! Sorry not to have responded earlier.

OpenStudy (mathmale):

Any loose ends to tie up? or are you all clear on this problem now?